# Self-interacting dark matter interpretation of Crater II

### v1

@@ Abstract
The satellite galaxy Crater II of the Milky Way is extremely cold and exceptionally diffuse. These unusual properties are challenging to understand in the standard model of cold dark matter. We investigate the formation of Crater II in self-interacting dark matter (SIDM), where dark matter particles can scatter and thermalize. We conduct a series of controlled N-body simulations to model the tidal evolution of Crater II, varying the self-interacting cross section, orbit parameters, and initial stellar distribution. Dark matter self-interactions lead to halo core formation and the distribution of stars expands accordingly. A cored SIDM halo also boosts tidal mass loss, allowing for a high orbit. Our simulations show that SIDM halos with a $1~{}{\rm kpc}$ core can simultaneously explain the low stellar velocity dispersion and the large half-light radius of Crater II, remaining robust to the initial distribution of stars. For the orbit motivated by the measurements from Gaia Early Data Release 3, the favored self-interacting cross section is approximately $60~{}{\rm cm^{2}/g}$ on the mass scale of Crater II.
I Introduction
@@
Over the past decade, astronomical surveys have discovered more than $30$ satellite dwarf galaxies associated with the Milky Way, see, e.g., Simon (2019). Among them, Crater II, located $117~{}{\rm kpc}$ away from the Sun, is particularly interesting, as it has the coldest velocity dispersion of $\sigma_{\rm los}\approx 2.7~{}{\rm km/s}$, while being exceptionally diffuse, with a projected half-light radius of $R_{1/2}\approx 1.07~{}{\rm kpc}$ Torrealba et al. (2016); Caldwell et al. (2017); Fu et al. (2019); Ji et al. (2021). Crater II has almost the lowest surface brightness among the satellites ever discovered. These unusual properties of Crater II make it an intriguing case for testing the standard model of cold dark matter (CDM).
@@
Environmental effects can be important for the formation of Crater II. The tides from its host Milky Way can strip away the halo mass and lower the velocity dispersion accordingly. Refs. Frings et al. (2017); Applebaum et al. (2021) conducted cosmological hydrodynamical CDM simulations and argued that Crater II-like systems could form through strong tidal interactions. However, more tailored studies Sanders et al. (2018); Borukhovetskaya et al. (2022) found that it is difficult to simultaneously realize both low velocity dispersion and large size of Crater II in CDM. The central cusp of a CDM halo is resilient to tidal stripping, and the tides must be sufficiently strong to explain the low velocity dispersion. Nevertheless, with such strong tides, the galaxy size would be truncated rapidly and it becomes much smaller than the observed size of Crater II; see Borukhovetskaya et al. (2022) for detailed discussion.
@@
In this work, we investigate the formation of Crater II within the scenario of self-interacting dark matter (SIDM). Dark matter self-interactions transport heat and thermalize the inner halo; see Tulin and Yu (2018); Adhikari et al. (2022) for reviews. For an SIDM halo in the core-expansion phase, a shallow density core forms and the stellar distribution expands accordingly. A cored halo also boosts tidal mass loss and mild tides are allowed for lowering the halo mass. We conduct controlled N-body simulations to model the evolution of Crater II in the tidal field of the Milky Way in SIDM, as well as in CDM for comparison. We will show that in SIDM the simulated galaxies can well reproduce the unusual properties of Crater II. There is a degeneracy between the tidal orbit and the self-interacting cross section. Taking the orbital parameters from Gaia Early Data Release 3 Gaia Collaboration (2021); McConnachie and Venn (2020), we will show that the required self-interacting cross section is $\sim 60~{}{\rm cm^{2}/g}$ for Crater II.
@@
The rest of the paper is organized as follows: In Sec. II, we introduce our simulation setup. In Sec. III, we discuss the properties of our simulated galaxies and compare them to the observations. We further discuss implications of our simulation results and conclude in Sec. IV.
II Simulation setup
@@
II.1 The dark matter halo of Crater II
We take three cases for the self-interacting cross section per particle mass: $\sigma/m=10~{}{\rm cm^{2}/g}$ (SIDM10), $30~{}{\rm cm^{2}/g}$ (SIDM30), and $60~{}{\rm cm^{2}/g}$ (SIDM60); as well as the CDM limit for comparison. For all cases, we model the initial halo of Crater II with a Navarro-Frenk-White (NFW) profile Navarro et al. (1997)
@@ II.1 The dark matter halo of Crater II
where $\rho_{s}$ and $r_{s}$ are the scale density and radius, respectively. For CDM, SIDM10, and SIDM30, we take $\rho_{s}=1.42\times 10^{7}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=2.06~{}{\rm kpc}$. The maximum circular velocity of the halo is $V_{\rm max}=26.57~{}{\rm km/s}$ and its associated radius is $r_{\rm max}=4.45~{}{\rm kpc}$. For SIDM60, $\rho_{s}=1.14\times 10^{7}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=2.23~{}{\rm kpc}$; $V_{\rm max}=25.77~{}{\rm km/s}$ and $r_{\rm max}=4.82~{}{\rm kpc}$. The halo parameters are overall consistent with those in Borukhovetskaya et al. (2022); Fattahi et al. (2018).
There are stringent constraints on the cross section on the scale of galaxy clusters $\sigma/m\lesssim 0.1~{}{\rm cm^{2}/g}$ for $V_{\rm max}\sim 1500~{}{\rm km/s}$ Kaplinghat et al. (2016); Harvey et al. (2015); Sagunski et al. (2021); Andrade et al. (2022), and a realistic SIDM model must have a velocity-dependent cross section that decreases towards massive halos. Nevertheless, we can use an effective constant cross section for a given halo to accurately characterize its gravothermal evolution Yang and Yu (2022); Outmezguine et al. (2023); Yang et al. (2023a, b); Fischer et al. (2023). Thus the constant cross sections we take should be regarded as effective cross sections for the Crater II halo.
We use the public code GADGET-2 Springel et al. (2001); Springel (2005), which is implemented with an SIDM module developed and tested in Yang and Yu (2022), and the code SpherIC Garrison-Kimmel et al. (2013) to generate the initial condition. For all cases, the initial halo mass is $3.37\times 10^{9}~{}{M_{\odot}}$ and there are $10^{7}$ simulation particles. The particle mass is $337~{}M_{\odot}$ and the Plummer softening length is $7~{}{\rm pc}$. The resolution is high enough for the purpose of this work.
II.2 The stellar component of Crater II
Since Crater II is dark matter-dominated, we can assume that stars are massless tracers of the halo potential and use the technique introduced in Bullock and Johnston (2005) to model the stellar component. For each simulation particle, we can attach an appropriate probability to represent a star. The equilibrium distribution function of stars is Errani and Peñarrubia (2020)
@@ II.2 The stellar component of Crater II
Ref. Torrealba et al. (2016) originally used a Plummer profile to describe the stellar distribution and reported the 2D half-light radius as $R_{1/2}=1.066~{}{\rm kpc}$. Ref. Borukhovetskaya et al. (2022) found that the Einasto profile in Equation 3 ($\alpha=1$) fits the Plummer projected density profile well. In this work, we take the same approach and choose the initial Einasto scale radius to be $r_{E}=0.40~{}{\rm kpc}$, $0.73~{}{\rm kpc}$, and $1.37~{}{\rm kpc}$ as in Borukhovetskaya et al. (2022). For later snapshots, we fit the stellar distribution using the Einasto profile as well, extract the corresponding $r_{E}$ value, and calculate the half-light radius as $R_{1/2}\approx 2.03r_{E}$.
II.3 The Milky Way
@@
    with $\rho_{s}=7.7\times 10^{6}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=20.2~{}{\rm kpc}$.
  • A spherical bulge of a Hernquist profile Hernquist (1990)
@@
  • Two axisymmetric disks of a Miyamoto-Nagai profile Miyamoto and Nagai (1975)
@@
These parameters of the Milky Way are the same as those in Borukhovetskaya et al. (2022); Errani and Peñarrubia (2020), which are motivated by the mass model in McMillan (2011) with a circular velocity of $239~{}{\rm km/s}$ at the solar radius $R_{\odot}=8.3~{}{\rm kpc}$. Since the main halo is treated as a static potential, we neglect the evaporation effect between the satellite halo and the main halo Nadler et al. (2020); Slone et al. (2023). The approximation is justified if the cross section in the main halo ($V_{\rm max}\approx 250~{}{\rm km/s}$) is below ${\cal O}(1)~{}{\rm cm^{2}/g}$. This condition can be satisfied naturally for velocity-dependent SIDM models that evade the cluster constraints, see, e.g., Yang et al. (2023c); Nadler et al. (2023) for examples.
II.4 Orbital parameters
Quantity Value Ref
$D_{\odot}$ [kpc] $117.5\pm 1.1$ Torrealba et al. (2016)
$v_{\rm los}$ [km/s] $87.5\pm 0.4$ Caldwell et al. (2017)
$\alpha_{\rm J2000}$ [deg] $117.3^{\circ}$ Torrealba et al. (2016)
$\delta_{\rm J2000}$ [deg] $-18.4^{\circ}$ Torrealba et al. (2016)
$\mu_{\alpha^{\star}}$ [mas/yr] $-0.246\pm 0.052$ Kallivayalil et al. (2018)
$-0.184\pm 0.061\pm 0.035$ Fritz et al. (2018)
$-0.07\pm 0.02$ McConnachie and Venn (2020)
$\mu_{\delta}$ [mas/yr] $-0.227\pm 0.026$ Kallivayalil et al. (2018)
$-0.106\pm 0.031\pm 0.035$ Fritz et al. (2018)
$-0.11\pm 0.01$ McConnachie and Venn (2020)
Table 1: Measured orbital parameters of Crater II: the distance from the Sun $D_{\odot}$, line-of-sight velocity in the solar rest frame $v_{\rm los}$, right ascension $\alpha_{\rm J2000}$ and declination $\delta_{\rm J2000}$, proper motions $\mu_{\alpha^{\star}}$ and $\mu_{\delta}$.
In Table 1, we summarize the measured orbital parameters of Crater II. Its proper motions are derived from data of the Gaia survey Brown et al. (2018, 2021) and there are quite large uncertainties among different analyses in the literature. We first consider three pairs of $\mu_{\alpha^{\star}}$ and $\mu_{\delta}$ as in Borukhovetskaya et al. (2022), denoted as as Orbits 1, 2, and 3, resulting in three pericenter distances in our simulation setup, $r_{\rm p}=2.5$, $13.8$, and $37.7~{}{\rm kpc}$, respectively; see Table 2. For Orbit 1, the pericenter in our setup is a factor of $1.7$ smaller than the one in Borukhovetskaya et al. (2022). For Orbits 2 and 3, the agreement is within $10\%$. We further consider five more orbits, denoted as Orbits 4-8, and their pericenter distance is evenly distributed within the range spanned by Orbits 2 and 3. The proper motions of Orbit 3 are well consistent with the measured values based on Gaia Early Data Release 3 McConnachie and Venn (2020).
[…]
O1 -0.169 -0.267 2.5 None
O2 -0.102 -0.225 13.8 ${}^{\dagger}$, 10
O3 -0.07 -0.11 37.7 CDM, 10, 30, ${}^{\dagger}$
O4 -0.097 -0.206 16.6 10, 30
O5 -0.091 -0.187 19.9 ${}^{\dagger}$, 30
O6 -0.086 -0.168 23.7 10, 30
O7 -0.0807 -0.148 27.9 10, ${}^{\dagger}$
O8 -0.0753 -0.129 32.6 30
[…]
III Results
@@
III.1 Tidal evolution of the halo mass
@@
Figure 1: Left panel: Evolution of the circular velocity at half-light radius $V(R_{1/2})\equiv\sqrt{GM(R_{1/2})/R_{1/2}}$ for the four benchmark cases: CDM (black), SIDM $\sigma/m=10~{}{\rm cm^{2}/g}$ (blue), $30~{}{\rm cm^{2}/g}$ (magenta), and $60~{}{\rm cm^{2}/g}$ (red), with their orbit parameters listed in Table 2. For comparison, the CDM case with Orbit 2 is also shown (green). The horizontal line and shaded region denote the measured value and its uncertainties $V(R_{1/2})=4.3\pm 0.5~{}{\rm km/s}$ Caldwell et al. (2017). The triangle symbol denotes the final snapshot for comparing the simulated galaxies with Crater II; they are relaxed after their last apocenter passage. Right panel: Corresponding circular velocity profiles of the benchmarks at the final snapshot. The gray curve denotes the initial condition. The data points with error bars represent the measurements $V(R_{1/2})$ and $V(\frac{4}{3}R_{1/2})$ from Caldwell et al. (2017).
In Figure 1 (left panel), we show the tidal evolution of the circular velocity at the half-light radius $V(R_{1/2})=\sqrt{GM(R_{1/2})/R_{1/2}}$, a proxy for the halo mass within $r=R_{1/2}=1.066~{}{\rm kpc}$, for the four benchmark cases: CDM (black) and SIDM10 (blue), SIDM30 (magenta), and SIDM60 (red), corresponding to Orbits 2, 5, 7 and 3, respectively. We also include the CDM case with Orbit 3 for comparison (green). The horizontal line and shaded region denotes the measured value and its uncertainties $V(R_{1/2})=4.3\pm 0.5~{}{\rm km/s}$ Caldwell et al. (2017). For our CDM case with Orbit 2, $V(R_{1/2})$ is close to the measured value after five pericenter passages. This is in good agreement with Borukhovetskaya et al. (2022), further validating our implementation.
For the three SIDM benchmarks, $V(R_{1/2})$ reaches the measured value after several pericenter passages, and a larger pericenter is required as the cross section increases. For Orbit 3, which is based on the proper motions from Gaia Early Data Release 3 McConnachie and Venn (2020), the pericenter is $r_{\rm p}=37.7~{}{\rm kpc}$ and an SIDM model with $\sigma/m=60~{}{\rm km/s}$ is needed, while it would be $V(R_{1/2})\approx 15~{}{\rm km/s}$ for a CDM halo. If we relax constraints on the orbital parameters, a smaller cross section works as well. For $\sigma/m=10~{}{\rm km/s}$ and $30~{}{\rm km/s}$, $r_{\rm p}=19.9~{}{\rm kpc}$ (Orbit 5) and ${27.9}~{}{\rm kpc}$ (Orbit 7), respectively. We choose the timescale marked by the triangle symbol as the final snapshot for analyzing other properties of the simulated galaxies, at which they are well relaxed. In Appendix, we will show that our main results are robust to this choice.
@@
Figure 1 (right panel) shows the corresponding circular velocity profiles at the final snapshot for the CDM and SIDM benchmarks, as well as the initial condition. The measurements at $r=R_{1/2}$ and $r=\frac{4}{3}R_{1/2}$ are $V(R_{1/2})=4.3^{+0.5}_{-0.5}~{}{\rm km/s}$ and $V(\frac{4}{3}R_{1/2})=4.8^{+0.6}_{-0.5}~{}{\rm km/s}$ Caldwell et al. (2017), respectively. Tidal stripping significantly reduces the overall halo mass, lowering the circular velocity in accord with the measurements at $r\sim R_{1/2}$ for both CDM and SIDM. However, their $V(r)$ profiles within $R_{1/2}$ are different. For the SIDM cases, $V(r)$ increases with $r$ linearly, indicating that the SIDM halos have a shallow density core in the inner region. The halo core size is comparable to $R_{1/2}\sim 1~{}{\rm kpc}$. For the CDM case, $V(r)$ increases sharply from the center and then decreases as its inner halo remains cuspy even after tidal stripping.
III.2 The stellar distribution

### v2

@@ Abstract
The satellite galaxy Crater II of the Milky Way is extremely cold and exceptionally diffuse. These unusual properties are challenging to understand in the standard model of cold dark matter. We use controlled N-body simulations to investigate the formation of Crater II in self-interacting dark matter (SIDM), where dark matter particles can scatter and thermalize. Taking the orbit motivated by the measurements from Gaia Early Data Release 3, we show a strong self-interacting cross section per particle mass of $60~{}{\rm cm^{2}/g}$ is favored for Crater II. The simulated SIDM halo, with a $1~{}{\rm kpc}$ core, leads to both a low stellar velocity dispersion and a large half-light radius for Crater II. These characteristics remain robust regardless of the initial stellar distribution.
1 Introduction
@@
Over the past decade, astronomical surveys have discovered more than $30$ satellite dwarf galaxies associated with the Milky Way, see, e.g., Simon (2019). Among them, Crater II, located $117~{}{\rm kpc}$ away from the Sun, is particularly interesting, as it has the coldest velocity dispersion of $\sigma_{\rm los}\approx 2.7~{}{\rm km/s}$, while being exceptionally diffuse, with a projected half-light radius of $R_{1/2}\approx 1.07~{}{\rm kpc}$ (Torrealba et al., 2016; Caldwell et al., 2017; Fu et al., 2019; Vivas et al., 2020; Ji et al., 2021). Crater II has almost the lowest surface brightness among the satellites ever discovered. These unusual properties of Crater II make it an intriguing case for testing the standard model of cold dark matter (CDM).
@@
Environmental effects can be important for the formation of Crater II. The tides from its host Milky Way can strip away the halo mass and lower the velocity dispersion. Frings et al. (2017); Applebaum et al. (2021) conducted cosmological hydrodynamical CDM simulations and argued that Crater II-like galaxies could form through strong tidal interactions. However, more tailored studies found that it is difficult to simultaneously realize both low velocity dispersion and large size of Crater II in CDM (Sanders et al., 2018; Borukhovetskaya et al., 2022). The central cusp of a CDM halo is resilient to tidal stripping, and the tides must be sufficiently strong to explain the low velocity dispersion. Nevertheless, with such strong tides, the galaxy size would be truncated rapidly and it becomes much smaller than the observed size of Crater II; see Borukhovetskaya et al. (2022) for detailed discussion. It remains to be seen if baryonic feedback can modify these CDM predictions in accord with the observations of Crater II.
@@
In this work, we investigate the formation of Crater II within the scenario of self-interacting dark matter (SIDM). Dark matter self-interactions transport heat and thermalize the inner halo; see Tulin & Yu (2018); Adhikari et al. (2022) for reviews. For an SIDM halo in the core-expansion phase, a shallow density core forms and the stellar distribution correlates with the core size. A cored halo also boosts tidal mass loss and mild tides are allowed for lowering the halo mass. We conduct controlled N-body simulations to model the evolution of Crater II in the tidal field of the Milky Way in SIDM, as well as in CDM for comparison. We will show that in SIDM the simulated galaxies can well reproduce the unusual properties of Crater II. Taking the orbital parameters from Gaia EDR3 (Gaia Collaboration, 2021; McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022), we will show that the required self-interacting cross section is $\sim 60~{}{\rm cm^{2}/g}$ for Crater II.
@@
The rest of the paper is organized as follows: In Section 2, we introduce our simulation setup. In Section 3, we discuss the properties of our simulated galaxies and compare them to the observations. We further discuss implications of our simulation results and conclude in Section 4. In Appendix A, we discuss the choice of the final snapshot. In Appendix B, we show the degeneracy effect between tidal orbit and cross section. In Appendix C, we present additional N-body simulations with live stellar particles and further confirm the results discussed in the main text.
2 Simulation setup
@@
2.1 The dark matter halo of Crater II
We consider three cases for the self-interacting cross section per particle mass: $\sigma/m=10~{}{\rm cm^{2}/g}$ (SIDM10), $30~{}{\rm cm^{2}/g}$ (SIDM30), and $60~{}{\rm cm^{2}/g}$ (SIDM60); as well as the CDM limit for comparison. For all cases, we model the initial halo of Crater II with a Navarro-Frenk-White (NFW) profile (Navarro et al., 1997)
@@ 2.1 The dark matter halo of Crater II
where $\rho_{s}$ and $r_{s}$ are the scale density and radius, respectively. For CDM, SIDM10, and SIDM30, we take $\rho_{s}=1.42\times 10^{7}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=2.06~{}{\rm kpc}$. The maximum circular velocity of the halo is $V_{\rm max}=26.57~{}{\rm km/s}$ and its associated radius is $r_{\rm max}=4.45~{}{\rm kpc}$. For SIDM60, $\rho_{s}=1.14\times 10^{7}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=2.23~{}{\rm kpc}$; $V_{\rm max}=25.77~{}{\rm km/s}$ and $r_{\rm max}=4.82~{}{\rm kpc}$. The halo parameters are overall consistent with those for progenitors of simulated satellite galaxies that have similar stellar masses of Crater II (Fattahi et al., 2018). As we will show later, with the tidal orbit from the measurements of Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022), only SIDM60 can reproduce the observations of Crater II. Thus we will focus on SIDM60 and CDM for the rest of the paper, and present SIDM10 and SIDM30 in Appendix B to show the degeneracy effect between orbit and cross section.
There are stringent constraints on the cross section on the scales of galaxy clusters $\sigma/m\lesssim 0.1~{}{\rm cm^{2}/g}$ for $V_{\rm max}\sim 1500~{}{\rm km/s}$ (Kaplinghat et al., 2016; Harvey et al., 2015; Sagunski et al., 2021; Andrade et al., 2022; Rocha et al., 2013; Peter et al., 2013) and elliptical galaxies (Kong et al., 2024). Thus, a viable SIDM model must have a velocity-dependent cross section, which is strong in dwarf galaxies, while diminishing towards the clusters; see Tulin & Yu (2018) and references therein for the discussion about particle physics realizations of SIDM. Nevertheless, we can use an effective constant cross section for a given halo to accurately characterize its gravothermal evolution (Yang & Yu, 2022; Outmezguine et al., 2023; Yang et al., 2023c, b; Fischer et al., 2023). The constant cross sections we take should be regarded as effective cross sections for the Crater II halo.
We use the public code GADGET-2 (Springel et al., 2001; Springel, 2005), which is implemented with an SIDM module developed and tested in Yang & Yu (2022), and the code SpherIC (Garrison-Kimmel et al., 2013) to generate the initial condition. For all cases, the initial halo mass is $3.37\times 10^{9}~{}{M_{\odot}}$ and there are $10^{7}$ simulation particles. The particle mass is $337~{}M_{\odot}$ and the Plummer softening length is $7~{}{\rm pc}$. The resolution is high enough for the purpose of this work.
2.2 The stellar component of Crater II
Since Crater II is dark matter-dominated, we can assume that stars are massless tracers of the halo potential and use the technique introduced in Bullock & Johnston (2005) to model the stellar component, i.e., the tagging method. In this work, we have further tested the method and validated its application in SIDM using N-body simulations with live stellar particles; see Appendix C.
For each simulation particle, we can attach an appropriate probability to represent a star. The equilibrium distribution function of stars is (Errani & Peñarrubia, 2020)
@@ 2.2 The stellar component of Crater II
Torrealba et al. (2016) originally used a Plummer profile to describe the stellar distribution and reported the 2D half-light radius as $R_{1/2}\approx 1.066~{}{\rm kpc}$Borukhovetskaya et al. (2022) found that the Einasto profile in Equation 3 ($\alpha=1$) fits the Plummer projected density profile well. For the purpose of comparison, we take the same approach and choose the initial Einasto scale radius to be $r_{E}=0.40~{}{\rm kpc}$, $0.73~{}{\rm kpc}$, and $1.37~{}{\rm kpc}$ as in Borukhovetskaya et al. (2022). Since the 2D half-light radius is $R_{1/2}=2.03r_{E}$, we have the initial values $R_{1/2}\approx 0.8~{}{\rm kpc}$, $1.5~{}{\rm kpc}$, and $2.8~{}{\rm kpc}$. These values bracket the measured value of $R_{1/2}\approx 1~{}{\rm kpc}$, and they are overall consistent with those found for progenitors of satellite galaxies that have similar stellar masses of Crater II in hydrodynamical CDM simulations (Fattahi et al., 2018).
For later snapshots, we fit the stellar distribution using the Einasto profile as well, extract the corresponding $r_{E}$ value, and calculate the half-light radius using the relation $R_{1/2}=2.03r_{E}$. We have checked that using a Plummer profile in the tagging method gives rise to a similar evolution trajectory of the stellar distribution as with the Einasto profile. Furthermore, we have also confirmed that with the initial stellar distributions assumed in this work, the dark matter mass completely dominates over the stellar component in the entire evolution history of the simulated halo and the tagging method is well justified.
2.3 The Milky Way
@@
    with $\rho_{s}=7.7\times 10^{6}~{}M_{\odot}/{\rm kpc}^{3}$ and $r_{s}=20.2~{}{\rm kpc}$. The corresponding halo mass is $M_{200}\approx 1.15\times 10^{12}~{}M_{\odot}$ and the maximum halo circular velocity is $V_{\rm max}\approx 192~{}{\rm km/s}$.
  • A spherical bulge of a Hernquist profile (Hernquist, 1990)
@@
  • Two axisymmetric disks of a Miyamoto-Nagai profile (Miyamoto & Nagai, 1975)
@@
These values are the same as those in Borukhovetskaya et al. (2022), which are motivated by the Milky Way mass model in McMillan (2011) with a circular velocity of $239~{}{\rm km/s}$ at the solar radius $R_{\odot}=8.3~{}{\rm kpc}$. We have checked that the pericenter of our simulated satellite halo would shift a few percent if taking the parameter values from the more recent model in McMillan (2016). The main halo is treated as a static potential, and we neglect the evaporation effect between the satellite halo and the main halo (Nadler et al., 2020; Slone et al., 2023). The approximation is justified if the cross section in the main halo is below ${\cal O}(1)~{}{\rm cm^{2}/g}$. This condition can be satisfied naturally for velocity-dependent SIDM models that evade the cluster constraints, see, e.g., Yang et al. (2023a); Nadler et al. (2023) for examples.
2.4 Orbital parameters
$D_{\odot}$ $117.5\pm 1.1$ Torrealba et al. (2016)
$v_{\rm los}$ $87.5\pm 0.4$ Caldwell et al. (2017)
$\alpha_{\rm J2000}$ $117.3^{\circ}$ Torrealba et al. (2016)
$\delta_{\rm J2000}$ $-18.4^{\circ}$ Torrealba et al. (2016)
$\mu_{\alpha^{\star}}$ $-0.07\pm 0.02$ McConnachie & Venn (2020)
$\mu_{\delta}$ $-0.11\pm 0.01$ McConnachie & Venn (2020)
Table 1: Measured orbital parameters of Crater II: the distance from the Sun $D_{\odot}$ [kpc], line-of-sight velocity in the solar rest frame $v_{\rm los}$ [km/s], right ascension $\alpha_{\rm J2000}$ and declination $\delta_{\rm J2000}$ [deg], proper motions $\mu_{\alpha^{\star}}$ and $\mu_{\delta}$ [mas/yr].
In Table 1, we summarize the measured orbital parameters of Crater II. Its proper motions are derived from data of the Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022). In our simulation we have tested a series of obits with a wide range of pericenter $r_{\rm p}\approx 2.5\textup{--}37.7~{}{\rm kpc}$. Our main results are based on two orbits, denoted as O2 and O3, as listed in Table 2, where we have used the same naming scheme as in Borukhovetskaya et al. (2022). The orbit O2 has a pericenter of $r_{\rm p}\approx 13.8~{}{\rm kpc}$ and it is designed such that the CDM halo can lose sufficient mass, while the orbit O3 has $r_{\rm p}\approx 37.7~{}{\rm kpc}$ and it is well consistent with the measurements of Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022).
Orbit $\mu_{\alpha^{\star}}$ $\mu_{\delta}$ $r_{\rm p}$ Model
O2 -0.102 -0.225 13.8 CDM
O3 -0.07 -0.11 37.7 CDM, SIDM60
Table 2: Orbital parameters considered for our CDM and SIDM simulations. Columns from left to right: Orbit name label, proper motions $\mu_{\alpha^{\star}}$ and $\mu_{\delta}$ [mas/yr], pericenter distance [kpc], and dark matter model. The orbit O3 is consistent with the measurements of Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022), while O2 is designed for the CDM halo to lose sufficient mass.
3 Results
@@
3.1 Tidal evolution of the halo mass
@@
Figure 1: Left panel: Evolution of the circular velocity at half-light radius $V(R_{1/2})\equiv\sqrt{GM(R_{1/2})/R_{1/2}}$ for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). The horizontal line and shaded region denote the measured value and its uncertainties $V(R_{1/2})=4.3\pm 0.5~{}{\rm km/s}$ (Caldwell et al., 2017). The triangle symbol denotes the final snapshot for comparing the simulated galaxies with Crater II; they are relaxed after their last pericenter passage. Right panel: Corresponding circular velocity profiles at the final snapshot. The gray curve denotes the initial condition. The data points with error bars represent the measurements $V(R_{1/2})$ and $V(\frac{4}{3}R_{1/2})$ from Caldwell et al. (2017). The orbit O3 is consistent with the measurements of Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022), while the orbit O2 is designed such that the CDM halo has $V(R_{1/2})\approx 5~{}{\rm km/s}$ at the final snapshot.
In Figure 1 (left panel), we show the tidal evolution of the circular velocity at the half-light radius $V(R_{1/2})=\sqrt{GM(R_{1/2})/R_{1/2}}$, a proxy for the halo mass within $r=R_{1/2}=1.066~{}{\rm kpc}$, for the three cases: SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). The horizontal line and shaded region denotes the measured value and its uncertainties $V(R_{1/2})=4.3\pm 0.5~{}{\rm km/s}$ (Caldwell et al., 2017). We choose the timescale marked by the triangle symbol as the final snapshot for analyzing other properties of the simulated galaxies, at which they are well relaxed. In Appendix A, we will show that our main results are robust to this choice.
For SIDM60 O3, $V(R_{1/2})$ reaches the measured value after four pericenter passages. Thus, with the orbit motivated by the Gaia EDR3 measurements, an SIDM model with $\sigma/m\approx 60~{}{\rm cm^{2}/g}$ is needed as the pericenter is large $r_{\rm p}=37.7~{}{\rm kpc}$. For CDM O3, $V(R_{1/2})\approx 15~{}{\rm km/s}$, a factor of $3$ larger than the measured value. If we choose the orbit O2, the CDM halo can reach $V(R_{1/2})\approx 5~{}{\rm km/s}$ after five pericenter passages. However, its pericenter $r_{\rm p}\approx 13.8~{}{\rm kpc}$ is a factor of $2.7$ smaller than that from the EDR3 measurements.
@@
Figure 1 (right panel) shows the corresponding circular velocity profiles at the final snapshot for the three cases, as well as the initial condition. The measurements at $r=R_{1/2}$ and $r=\frac{4}{3}R_{1/2}$ are $V(R_{1/2})=4.3^{+0.5}_{-0.5}~{}{\rm km/s}$ and $V(\frac{4}{3}R_{1/2})=4.8^{+0.6}_{-0.5}~{}{\rm km/s}$ (Caldwell et al., 2017), respectively. Tidal stripping can significantly reduce the overall halo mass, lowering the circular velocity in accord with the measurements at $r\sim R_{1/2}$ for SIDM60 O3 and CDM O2. However, their $V(r)$ profiles are different, which can be understood as follows. For CDM O2, $r_{\rm max}$ reduces to $0.4~{}{\rm kpc}$ at the final snapshot and the inner halo remains cusp. In the case of SIDM, the self-interactions push dark matter particles outwards. Thus, compared to CDM O2, SIDM60 O3 has a lower inner density while larger $r_{\rm max}\approx 1.5~{}{\rm kpc}$ at the final snapshot. We also see that the circular velocity of CDM O3 is too high to be consistent with the measurements.
3.2 The stellar distribution